### v1

# Robust Pronoun Use Fidelity with English LLMs: Are they Reasoning, Repeating, or Just Biased?

Abstract
Robust, faithful and harm-free pronoun use for individuals is an important goal for language models as their use increases, but prior work tends to study only one or two of these components at a time. To measure progress towards the combined goal, we introduce the task of pronoun use fidelity: given a context introducing a co-referring entity and pronoun, the task is to reuse the correct pronoun later, independent of potential distractors. We present a carefully-designed dataset of over 5 million instances to evaluate pronoun use fidelity in English, and we use it to evaluate 37 popular large language models across architectures (encoder-only, decoder-only and encoder-decoder) and scales (11M-70B parameters). We find that while models can mostly faithfully reuse previously-specified pronouns in the presence of no distractors, they are significantly worse at processing she/her/her, singular they and neopronouns. Additionally, models are not robustly faithful to pronouns, as they are easily distracted. With even one additional sentence containing a distractor pronoun, accuracy drops on average by 34%. With 5 distractor sentences, accuracy drops by 52% for decoder-only models and 13% for encoder-only models. We show that widely-used large language models are still brittle, with large gaps in reasoning and in processing different pronouns in a setting that is very simple for humans, and we encourage researchers in bias and reasoning to bridge them.
1 Introduction
Figure 1: We evaluate model accuracy at using the correct pronoun for an entity when provided with an explicit introduction and 0-5 distractor sentences. Llama-2-70B and RoBERTa-large show large accuracy drops with just one distractor.
Third-person pronouns (he, she, they, etc.) are words that construct individuals’ identities in conversations (Silverstein, 1985). In English, these pronouns mark referential gender for the entity they are referring to, which can also index an individual’s social gender, e.g., man, woman, non-binary (Cao and Daumé III, 2020). Correctly using the pronouns an individual identifies with is important, as misgendering (including through incorrect pronoun use) can in the best case be a social faux pas (Stryker, 2017) and in the worst case, cause psychological distress, particularly to transgender individuals (McLemore, 2018).
Accordingly, language technology should acknowledge the sensitivity of faithful and correct pronoun use. To this end, many studies have explored how large language models handle pronouns, showing that they stereotypically associate pronouns and occupations (Kurita et al., 2019), reason about co-referring pronouns and entities better when they conform to stereotypes (Tal et al., 2022), fail when exposed to novel pronoun phenomena such as neopronouns (Lauscher et al., 2023), and cannot consistently reuse neopronouns during generation (Ovalle et al., 2023a). These shortcomings create quality of service differences and cause representational harm, amplifying discrimination against certain pronoun users (Blodgett et al., 2020; Dev et al., 2021).
However, (i) studies on stereotypical associations tend to test for intrinsic bias, with brittle methodology (Seshadri et al., 2022) that may not translate to extrinsic bias or harms in downstream applications (Goldfarb-Tarrant et al., 2021), (ii) studies on reasoning with pronouns tend to focus on coreference resolution, the ability to reason about connections between pronouns and entities, which may not translate to a model’s use of pronouns, and (iii) studies on pronoun use tend to examine simplistic pronoun reuse with no more than one entity. A key research question that has gone unanswered thus far is: How robustly faithful are models to one’s pronouns? We address this research gap by presenting the first study measuring the robustness and fidelity of models’ pronoun use.
Contributions.(1) To investigate whether models can reason robustly about pronouns in context, we propose a new task, pronoun use fidelity: given a context in which a co-referring entity and pronoun are introduced, the task is to reconstruct the pronoun later, independent of potential distractors. (2) We present a novel, large-scale dataset consisting of 7,000 up to 2 million instances depending on the setup, carefully designed to evaluate this task. (3) We study pronoun use fidelity across 37 popular language models covering a range of architectures and scales, and we analyze model errors to examine whether models are reasoning, repeating, or just biased.
First, we examine pronoun predictions in the absence of context as previous studies have done, to establish a bias baseline and examine the consistency of model predictions across case (§5). Next, we evaluate whether models can overcome their biased pronoun predictions by showing them an introductory context to explicitly establish what pronoun to use (§6). We find that all models are good at this task, and bigger models are better. Then, we attempt to distract models by inserting additional sentences using a different pronoun to talk about another person (§7). Models are not robust to distractors, and even one distractor sentence vastly deteriorates model performance as shown in Figure 1. This result holds even at the scale of 70 billion parameter models, despite how simple this task is for humans. Finally, in a detailed error analysis (§8), we examine the extent to which model errors can be attributed to distraction or falling back to intrinsic bias. Overall, we show a large gap in robust model reasoning about pronouns in a simple setting, and we find that encoder-only and decoder-only models behave in fundamentally different ways, not only in their performance on this task, but also in the types of errors they make. We release all code and data at https://github.com/uds-lsv/pronoun-use-fidelity.
2 Pronoun Use Fidelity Task
Given a context in which a co-referring entity and pronoun are introduced, the task is to reconstruct the pronoun later in a sentence about the entity, independent of potential distractors.
[…]
Introduction: The accountant had just eaten a big meal so her stomach was full.
Distractor 1: The taxpayer needed coffee because their day had started very early.
Distractor N: Their sleep had been fitful.
Task sentence: The accountant was asked about ___ charges for preparing tax returns.
[…]
the probability $P$ of reconstructing the correct pronoun $p_{a}$ in the sentence $t({\color[rgb]{0.13,0.55,0.13}\definecolor[named]{pgfstrokecolor}{rgb}{0.13,0.55,0.13}e_{a}},p)$, given the context.
3 Data
We begin by discussing our dataset construction procedure to evaluate pronoun use fidelity at scale. It consists of a three-step pipeline allowing us to create naturalistic instances of our task while ensuring a correct answer: template creation (§3.1), template assembly (§3.2), and validation (§3.3).
Our data covers 60 occupations (see Appendix A) and four third-person pronouns (he, she, they and xe) in three cases (nominative, accusative and possessive dependent). Our occupations (and participants for the distractor sentences) are chosen from Winogender schemas (Rudinger et al., 2018), as their bias characteristics are well-studied in NLP fairness literature. In addition to the English standard masculine (he/him/his) and femininine (she/her/her) pronouns, we heed Lauscher et al.’s (2022) call for more inclusive NLP research by examining two more pronoun sets that are less well-studied in the fairness literature: singular they (they/them/their), the pronoun of choice of over 75% of respondents to the Gender Census (Lodge, 2023), and xe/xem/xyr, the most popular neopronoun according to the same census.
3.1 Template creation
Task templates. We create one task sentence template per occupation and pronoun case, for a total of 180 templates, designed to create an unambiguous coreference with the occupation only. For instance, charges for preparing tax returns can only belong to an accountant, never a taxpayer, which is the corresponding participant.
Context templates. Semantically bleached templates such as He is an accountant are well-established in the literature for testing word embedding associations (Caliskan et al., 2017), but it is unnatural to use more than one consecutively. On the other hand, natural corpora like Levy et al. (2021) contain occupation-specific sentences that cannot be used more generally. As a compromise between a more controlled setting and naturalistic data, we create templates with generic themes that apply to all humans, e.g., universal emotions and sensations (hungry/full, tired/energetic, unhappy/happy, etc.). We create 10 context templates per pronoun case, for a total of 30 templates.
3.2 Template assembly
We then combine and instantiate the previously created templates to assemble our data instances. First, we select a task sentence for a certain occupation and pick one of four pronouns in the correct case to use as the ground truth label for the task. The occupation becomes entity $e_{a}$ and the pronoun is $p_{a}$. We then create an introduction sentence by sampling one of the 10 context templates with the matching case, and instantiating it with the selected occupation $e_{a}$ and pronoun $p_{a}$. The simplest version of the pronoun use fidelity task includes just this introduction sentence followed by the task sentence. Instantiating 10 introductory templates with 4 different pronouns and pairing them with task templates for 60 occupations across 3 cases gives us a total of 7,200 unique instances for this version of the task.
To create more complex data instances, we insert a variable number of distractor sentences between the introduction and task sentences. The first distractor sentence always explicitly establishes a coreference between the participant and the distractor pronoun, and any subsequent distractor sentences implicitly continue this coreference by using just the pronoun. We first select a distractor pronoun $p_{b}$ from the three unused pronouns. We sample templates from the remaining context templates [^1] For more details on this sampling, refer to Appendix B., instantiating them with this pronoun and a distractor entity $e_{b}$, the participant that the occupation is paired with. For the one-distractor case, instantiating 4 templates with 3 unused pronouns gives us 86,400 unique instances.
Data type Number of instances
With no context
Task sentences 180
With introductory context
+ 0 distractors 3 x 2,160 0,00(of 7,200)
+ 1 distractors 3 x 2,160 0,0(of 86,400)
+ 2 distractors 3 x 2,160 0,(of 345,600)
+ 3 distractors 3 x 2,160 (of 1,036,800)
+ 4 distractors 3 x 2,160 (of 2,073,600)
+ 5 distractors 3 x 2,160 (of 2,073,600)
Table 1: Number of dataset instances. Pronoun use fidelity instances consist of task instances combined with introductory contexts and optional distractors. We subsample 3 sets of 2,160 sentences (of the total number of instances we created).
Our stackable dataset design allows us to generate a vast amount of data of varying lengths, which we believe reflects varying levels of difficulty for models. We subsample these datasets with three random seeds for the rest of our evaluation, ensuring that all occupations, cases, pronoun declensions and distractor pronouns are equally represented in each subsampled set of 2,160 sentences. Data statistics are shown in Table 1.
3.3 Data validation
To verify that the pronoun use fidelity task is valid and easy for humans, we validate a subset of the combined sentences, in addition to validating all the individual sentence templates. All annotator instructions are provided in Appendix C.
Sentence templates. Two authors with linguistic training iteratively created and validated the sentence templates for grammaticality and coreference correctness until we reached consensus. These were independently verified by an annotator who also rated 100% of the sentences as grammatical and with the correct coreferences.
Pronoun use fidelity task. We sampled 100 instances with each possible number of distractors (0-5), for a total of 600 instances. One author and one annotator had to fill in the pronoun and they each performed with 99.8% accuracy. [^2] They disagreed on non-overlapping instances which appeared to be random slips.
Model Sizes Architecture
Evaluated with (Pseudo) Log Likelihoods
ALBERT-v2 base (11M), large (17M), xlarge (58M), xxlarge (223M) Encoder-only
BERT base (110M), large (340M) Encoder-only
RoBERTa base (125M), large (355M) Encoder-only
MosaicBERT 137M Encoder-only
OPT 125M, 350M, 1.3B, 2.7B, 6.7B, 13B, 30B, 66B Decoder-only
Pythia 14M, 70M, 160M, 410M, 1B, 1.4B, 2.8B, 6.9B, 12B Decoder-only
Llama-2 7B, 13B, 70B Decoder-only
Evaluated with prompting
FLAN-T5 small (77M), base (248M), large (783), xl (2.85B), xxl (11.3B) Encoder-decoder
Llama-2-chat 7B, 13B, 70B Decoder-only
Table 2: Models we experiment with across a range of sizes and architectures.
4 Experimental Setup
We first list the models we experiment with, and explain our evaluation methods and metrics. Further details are provided in Appendix D.
4.1 Models
We experiment with a total of 37 transformer-based language models, chosen to evaluate the effects of architecture and scaling. Table 2 shows all models we experiment with. Our 9 encoder-only models are from the BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), ALBERT-v2 (Lan et al., 2020) and MosaicBERT (Portes et al., 2023) model families, as the first three remain well-used in NLP fairness literature, and the last is trained on much more data. As for our 20 decoder-only models, we select the popular Llama-2 (Touvron et al., 2023) model family, as well as OPT (Zhang et al., 2022) and Pythia (Biderman et al., 2023) for their large range of model sizes. We also experiment with eight popular chat models that are further trained with instruction-tuning and reinforcement learning, to evaluate their instruction-following behaviour on this task; specifically, we use decoder-only Llama-2-chat models (Touvron et al., 2023) and encoder-decoder FLAN-T5 models (Chung et al., 2022).
4.2 Obtaining predictions
Decoder-only models. For the majority of our experiments, we follow Hu and Levy (2023) in taking direct measurements of probabilities as a proxy for models’ metalinguistic judgements. We verbalize four versions of each data instance, i.e., we fill in the blank with each of the four pronouns we consider, creating four options. We then compute sentence-level model log likelihoods for these options, and select the option with the highest log likelihood as the model’s choice.
Encoder-only models. For comparable evaluation across model architectures and pronouns, we use pseudo log likelihoods for encoder-only models (Salazar et al., 2020; Kauf and Ivanova, 2023). We do not use masked token prediction due to tokenization issues with neopronouns (Ovalle et al., 2023b); briefly, we want xe to be tokenized as it normally would be (which is often as two tokens) rather than as a single UNK token.
Chat models. In line with their most naturalistic usage scenarios, we evaluate chat models (FLAN-T5 and Llama-2-chat) separately using prompting. We report the range of performance with 10 different prompts using boxplots, following Sclar et al. (2024). Our prompts, shown in Appendix E, were inspired by the prompts to elicit coreferences in the FLAN collection (Longpre et al., 2023).
4.3 Metrics
On the task sentences with no context, as there is no correct answer, we calculate the Pearson correlation coefficient between each model’s nominative, accusative and possessive pronoun predictions to evaluate the models’ case consistency.
On the pronoun fidelity task, as there is a correct answer, we compute accuracy. We report mean accuracy over the three randomly sampled subsets of our dataset, and show the standard deviation with error bars or shading. Where possible, we perform significance testing with a Welch’s t-test and a threshold of 0.05. We use human performance as our ceiling, and compare models evaluated with (pseudo) log likelihoods to a baseline of randomly selecting 1 of the 4 pronouns (i.e., 25%).
5 Model Predictions with No Context
Model pronoun prediction with no context is the task studied in numerous papers on pronominal gender bias (Kurita et al., 2019, inter alia), which show associations between occupations and pronouns based on social gender stereotypes, e.g., doctor-he and nurse-she. However, most of them ignore cases beyond the nominative (Munro and Morrison, 2020). If model associations between occupations and pronouns are primarily explained by social bias and occupational statistics, then we should see correlations between pronoun predictions across cases. In other words, if a model associates accountants with he, then it might also associate accountants with him and his. Therefore, we first examine model pronoun predictions on our task sentences with no context, and evaluate the case consistency of pronoun predictions.
Figure 2: Distribution of Pearson correlations of per-occupation pronoun predictions of different cases: nom(inative), acc(usative) and poss(essive).
Figure 3: Pronoun use fidelity of all models with an introductory context and no distractors. Accuracy is averaged across all pronouns and cases, and is above chance (0.25) but below human performance (1.0).
he/him/his Refer to caption
she/her/her Refer to caption
they/them/their Refer to caption
xe/xem/xyr Refer to caption
Figure 4: Pronoun use fidelity with an introductory context and no distractors, split by pronoun series.
Figure 2 shows the distribution of Pearson correlations between per-occupation pronoun predictions of different cases (individual model correlations are provided in Appendix F). The range is large, e.g., --0.46 for RoBERTa-base compared to 0.47 for Pythia-160M for nominative-accusative prediction correlations, showing that there is a lot of variation in pronoun associations across case and models. Only nominative and possessive pronoun predictions are positively correlated for all but one model (ALBERT-base-v2), and accusative pronouns seem to behave quite differently; Llama-2-13B, for instance, has the highest nominative-possessive correlation of all models, 0.55, but nominative-accusative and accusative-possessive correlations of 0.00 and --0.01 respectively. As pronoun predictions seem not to line up well across case, these results suggest that biased pronoun predictions in this context just mean statistically biased predictions, adding to the existing evidence that templates are a brittle method of evaluating social biases in models (Selvam et al., 2023; Seshadri et al., 2022). Next, we examine whether models can override their intrinsic statistical biases on these templates when provided with a pronoun to use.
6 Injecting an Introductory Context
When models are provided with an introductory sentence explicitly establishing the pronoun to use for an entity, can they use that pronoun to refer to the same entity in the immediate next sentence?
[…]
Example: The accountant had just eaten a big meal so her stomach was full. The accountant was asked about ___ charges for preparing tax returns.
[…]
As Figure 3 shows, all models can use a simple introduction better than chance (up to 0.95 with MosaicBERT), but not as well as humans, who achieve perfect performance. We also see improvements with increasing model scale, with the exception of ALBERT-v2. Model accuracy on accusative pronouns is statistically significantly higher than both possessive and nominative pronouns, but patterns across models and scaling are the same.
(a) With 0 distractors
(b) With 5 distractors
Figure 5: Scaling behaviour by architecture. With 0 distractors (above), encoder-only models are comparable to decoder-only models that are orders of magnitude larger. With 5 distractors (below), encoder-only models are far better than all decoder-only models.
Which pronouns are harder? Even in the simplest case of the pronoun use fidelity task, patterns emerge when split by pronoun, as shown in Figure 4. Overall model accuracy on he/him/his is significantly higher than she/her/her, which in turn is significantly higher than both they/them/their and xe/xem/xyr, in line with previous findings that language technology has gaps when it comes to gender-neutral pronouns, especially neopronouns (Lauscher et al., 2023). Models show intriguing patterns with these last two pronoun sets. Most encoder-only models appear to handle the neopronoun better than singular they (e.g., BERT-large has an accuracy of 0.78 on xe/xem/xyr compared to 0.60 on they/them/their), which warrants further investigation. Decoder-only models smaller than 6.7B parameters struggle with the neopronoun, with every OPT and Pythia model smaller than 2.7B parameters performing below chance, and in some cases (e.g., Pythia-14M, Pythia-70M and Pythia-160M) even performing close to 0.0. Beyond this scale, however, models perform better on xe/xem/xyr than on singular they, with Llama-13B achieving 0.96 accuracy on the neopronoun. These differences are statistically significant. As the training data for individual model families is the same, this might suggest that decoder-only models could learn to generalize to novel pronouns starting at the scale of 6.7B parameters. This could also explain the poor performance that some previous studies of neopronouns find, as the largest model that Hossain et al. (2023) experiments with, for instance, is OPT-6.7B. The lower performance of bigger models with singular they could also be a reflection of human processing difficulties with definite, specific (singular) they, as has been observed in linguistics (Conrod, 2019).
7 Adding Distractors
To further probe whether models actually use “reasoning” when supplied with an introductory context, we systematically inject sentences containing distractor pronouns between the introduction and the task, to examine how this affects performance.
[…]
Example: The accountant had just eaten a big meal so her stomach was full. The taxpayer needed coffee because their day had started very early. Their sleep had been fitful. The accountant was asked about ___ charges for preparing tax returns.
Correct answer: her
[…]
Figure 5 compares results on the easiest case of pronoun use fidelity (no distractors) with the hardest case (5 distractors), split by architecture. Surprisingly, with no distractors, encoder-only models are much better than decoder-only models of the same scale, and their performance is comparable to or better than decoder-only models that are orders of magnitude larger; RoBERTa-base (125M) is 0.86 accurate compared to OPT-125M’s 0.55, and exceeds OPT-66B’s 0.83 despite being more than 500 times smaller. In the hardest version of our task with five distractors, encoder-only models are far better than all decoder-only models, which show dramatically degraded performance. Llama-70B only achieves 0.37 accuracy, compared to MosaicBERT’s impressive 0.87. The lack of robustness of decoder-only models to distractors is striking, given that most state-of-the-art models in popular use today are decoder-only models.
(a) OPT
(b) Pythia
(c) Llama-2
(d) ALBERT-v2
(e) BERT
(f) RoBERTa
Figure 6: Model behaviour with additional distractors. Decoder-only models (above) get steadily worse, whereas encoder-only models (below) get worse with one distractor and then slowly recover, plateauing below their no-distractor performance.
Examining these results granularly, we see more interesting patterns across architectures and scales, shown in Figure 6. All decoder-only models get steadily worse as distractors are added, whereas encoder-only models perform the worst with one distractor and then seem to slowly recover, never quite reaching their level of performance with no distractors. Scaling seems to hold within model families, with larger models performing better with more distractors than smaller models of the same type.
Our results highlight the need to investigate task performance on a variety of architectures, as they might not pattern similarly. We hypothesize that the vastly different behaviour of encoder-only and decoder-only models with additional distractors reflects their different pre-training goals: encoder-only models are trained to learn rich representations of text, whereas decoder-only models are trained to predict the next token, which might make them more prone to recency bias.
Which cases and pronouns are harder with distractor sentences? With five distractor sentences, model accuracy with different pronoun sets follows the same statistically significant patterns as before. Interestingly, accusative pronoun performance is significantly different from the other cases as before, but in the opposite direction, i.e., worse than performance on nominative and possessive pronouns. This change from the best-performing pronoun case to the worst could be explained by a tendency for models to simply repeat the most recent accusative pronoun, a hypothesis we test next by examining model distractibility.
8 Distractibility versus Bias
In this section, we focus on the error cases for an in-depth analysis to confirm that models make mistakes because they are distracted by our distractor sentences. When a model gets the answer wrong, it is for one of three reasons: (1) distractibility, i.e., repeating the distractor pronoun, (2) bias, i.e., reverting to the model’s context-free prediction, or (3) picking a completely different pronoun. All three possibilities are illustrated below, and we hypothesize that the first two possibilities are much more frequent than the third.
In cases where the distractor pronoun is the same as the model’s context-free prediction, it is impossible to disentangle distractibility and bias just from the model’s prediction. Hence, we exclude these and focus on the unambiguous error cases. We find that 74-93% of unambiguous model errors can be attributed to the interplay of model distractibility and bias, confirming our hypothesis.
Refer to captionRefer to caption
(a) OPT
(b) Pythia
(c) Llama-2
(d) ALBERT-v2
(e) BERT
(f) RoBERTa
Figure 7: Trends in model distractibility (use of the distractor pronoun) and model bias (reverting to the context-free prediction). With more distractors, the proportion of errors due to distraction increases for decoder-only models (above) and decreases for encoder-only models (below).
[…]
Context-free
Example: The accountant was asked about ___ charges for preparing tax returns.
Prediction: his
With introduction and distractors
Example: The accountant had just eaten a big meal so her stomach was full. The taxpayer needed coffee because their day had started very early. Their sleep had been fitful. The accountant was asked about ___ charges for preparing tax returns.
Correct answer: her
Distraction error: their
Bias error: his
Other error: xyr
[…]
Are errors with distractor sentences actually due to distraction? We first examine model distractibility, i.e., what percentage of errors are caused by models repeating the distractor pronoun instead of the correct pronoun. As expected, Figure 7 shows that across models, distraction is indeed the primary type of error for most models. Decoder-only models get increasingly distractible with more distractors, i.e., the proportion of errors due to distractor pronoun repetition steadily increases as distractors are added, saturating just below 85%. On the other hand, encoder-only models seem to become less distractible with the addition of more distractors. We know from the previous section that encoder-only models recover in their pronoun use fidelity with 2-5 distractors, but here we measure distractibility as a percentage of all errors. Thus, a constant or increasing proportion of all the model errors could be due to distraction, and the fact that it isn’t for encoder-only models is quite surprising! We leave it to future work to investigate whether this behaviour is related to positional bias or context length.
When examining distractibility by pronoun case, we find that model distractibility is consistently higher with accusative pronouns compared to other cases. In combination, the high model performance on accusative pronouns with no distractors and the low performance with distractors strongly suggests that models tend to simply repeat the most recent pronoun, rather than reverting to biased predictions or using anything resembling robust “reasoning.”
As encoder-only models make proportionally fewer distraction errors, their errors due to bias go up. With BERT-large in particular, as soon as there is more than one distractor, the biggest proportion of errors is due to bias rather than distraction. BERT-large appears more biased and less distractible than BERT-base, in contrast to all other models. Generally, larger models seem to be more distractible and revert to their bias less often, whereas smaller models are more biased and less distractible. Our findings on bias errors contrast with Tal et al. (2022), where larger models make a higher proportion of bias errors on a downstream task than smaller models. We hypothesize the differences in our results could be explained by the presence of distractors in our task, which clearly have a very strong influence on model behaviour in this setting.
9 Prompting Shows Similar Results
Prompting is a different mechanism compared to evaluating log likelihoods and could in theory produce very different results. In practice, however, standard prompting has the same patterns of performance as before with an additional set of models, as shown in Figure 8.
(a) FLAN-T5
(b) Llama-2-chat
Figure 8: Performance of chat models with additional distractors, using prompting.
Bigger models are mostly better and degrade more gracefully than the smaller ones, but there remains a lot of variance across prompts, as shown in the box plots. We note that for Llama-2-chat, the best results with zero distractors are worse than the Llama-2 results with log likelihood evaluation. However, prompting seems to perform better with the addition of distractors, suggesting that chain-of-thought prompting—although excessive for a task this simple—might be a way of boosting performance.
10 Discussion and Future Work
Our results show that even the biggest models of today are not up to the task of pronoun use fidelity once it includes a single distractor sentence. All models are easily distracted, but encoder-only models and decoder-only models show very different patterns both in performance degradation with more distractors and their reasons for errors. Beyond architecture, model predictions across pronoun cases even in the absence of context exhibit a lot of variation, and performance on this type of reasoning task should be evaluated carefully, with attention to how the overall patterns break down by pronoun and case. Below we expand on some questions raised by our findings.
“Reasoning” and “bias.” Throughout the paper, we refer to “reasoning” and “bias” but these are inaccurate terms, both as we use them and as NLP as a field does. What looks like “reasoning” when we inject an introductory context starts to look much more like repetition—or stochastic parroting (Bender et al., 2021)—when we add a distractor and see the same models performing drastically worse. Even the higher performance of encoder-only models cannot accurately be attributed to “reasoning” in the same way that we use this word for humans, as these models are not grounded in meaning from the real world (Bender and Koller, 2020). As these are all language models, it is more accurate to say that the way that decoder-only models model language is prone to repetition of recent examples of the same word class, compared to encoder-only models. As for “bias,” we show that statistical bias does not always overlap in predictable ways with social bias, by finding poor correlations across cases for model predictions. This has implications for NLP literature on model bias, where using sentence templates and nominative pronouns to quantify social bias in models is a popular methodology.
Why exactly do we see the patterns we see? Our dataset design and error analysis show us how these models behave differently on the same task with a comparable evaluation method, and allow us to disentangle the effects of repetition, distraction and statistical bias. However, it is beyond the scope of this paper to investigate where in the model or data this comes from and what we can do about it. Tools from interpretability literature, e.g., attribution analysis, could help here, and are an important direction for future work.
What can we learn from humans? The author who validated the pronoun use fidelity task declared it easy and unambiguous, regardless of the number of distracting sentences. Clearly models do not correctly assign likelihood to the same extent, and prompting performance seems limited as well. One of the reasons for human success at this task might be our ability to assign pronouns to people and track them as separate individuals. Chain-of-thought prompting, or a different way of encouraging the model to explicitly track people and their pronouns, might therefore be a way of boosting model performance on this task.
Beyond our dataset. Given the breadth of our task definition, future work could include examining pronoun use fidelity for participants, for names by extending Hossain et al. (2023), with differently ordered sentences, and in less artificial settings such as stories (Mostafazadeh et al., 2016) or with naturally collected data as in Webster et al. (2018) and Levy et al. (2021).
11 Related Work
The main areas of overlap with our work are: reasoning with pronouns, pronouns and occupational bias, and the effect of context on model predictions. The three studies that are closest to our work are Hossain et al. (2023) and Ovalle et al. (2023a), both of which explore pronoun reuse with an explicitly established coreference between an individual and a pronoun series, as well as Sharma et al. (2022), which injects context with an explicit coreference to encourage faithful pronoun translation. Unlike our work, however, none of these papers explore robustness via the addition of adversarial distractors.
Reasoning with pronouns. Most existing work about language model reasoning with pronouns focuses on the task of coreference resolution, typically using Winograd schemas (Levesque et al., 2012; Emelin and Sennrich, 2021). Of closest relevance to our work is Abdou et al. (2020), which shows that humans perform better than LLMs on perturbed versions of Winograd schemas designed to test generalization. Although Winograd schemas traditionally include reasoning about inanimate objects or groups, Winograd-like schemas that focus on third person singular pronouns and people entities have also been explored. The antecedents are sometimes named (Webster et al., 2018; Zhao et al., 2018) or referred to with a noun phrase, e.g., an occupation (Rudinger et al., 2018; Levy et al., 2021). Most studies focus on he and she, but recent work has expanded to include evaluation with singular they (Baumler and Rudinger, 2022) and with neopronouns (Cao and Daumé III, 2021; Felkner et al., 2023). Nevertheless, studies on coreference resolution evaluate whether systems can reason to identify the connection between a pronoun and an entity, without measuring whether LLMs can model faithful reuse of that pronoun later, as in our work.
Pronouns and occupational bias. Associations between pronouns and occupations have been studied extensively in NLP, in the subfields of fairness and machine translation. Many fairness studies examine intrinsic bias in masked token prediction (Kurita et al., 2019; de Vassimon Manela et al., 2021; Tal et al., 2022) in embeddings (Bolukbasi et al., 2016; Zhao et al., 2019; Gonen and Goldberg, 2019), and in translations (Stanovsky et al., 2019; Saunders et al., 2020; Iluz et al., 2023; Ghosh and Caliskan, 2023), particularly in context-aware machine translation (Müller et al., 2018; Voita et al., 2018; Fernandes et al., 2023). Work on translation bias with unambiguous sources and targets is closely connected to our English-only work, as there is a ground truth and therefore a way to measure accuracy and extrinsic bias.
The effect of context. The impact of context on the robustness of language model reasoning has been investigated in many areas other than pronoun fidelity, e.g., negation (Gubelmann and Handschuh, 2022), linguistic acceptability (Sinha et al., 2023), natural language inference (Srikanth and Rudinger, 2022), and question answering  (Liu et al., 2024; Levy et al., 2024).
12 Conclusion
We introduce the task of pronoun use fidelity and a corresponding dataset we designed to evaluate robust, faithful and harm-free pronoun use in language models, while disentangling the effects of distraction, repetition and bias. We find evidence of faithful pronoun reuse only in a very simple setting, where models are provided with an introductory context and no distractor pronouns. Even here, models show significant performance disparities with neopronouns, singular they and she/her/her, compared to he/him/his. Models are also not robust to more complex settings; even adding a single sentence containing a distractor pronoun, accuracy drops dramatically. Interestingly, encoder-only models are better at faithfully reusing pronouns than decoder-only models of the same scale, and they are also more robust to the addition of distractor sentences. We show that most model errors are indeed due to distraction, but with additional distractor sentences, encoder-only models become less distracted, while decoder-only models get even more distracted. Our results show that in a very simple setting for humans, widely-used large language models are unable to robustly and faithfully reason about pronouns, and continue to amplify discrimination against users of certain pronouns. We encourage researchers in bias and reasoning to bridge the performance gaps we report.
Limitations
Our task as it is defined in Section 2 is much broader than the scope of our dataset. We focus on occupations due to the wide attention they have received in prior literature, but we continue a long tradition of ignoring biases relating to the participants, e.g., child, taxpayer, etc. Additionally, we evaluate on a very easy version of this task that allows us to quantify repetition, i.e., the case of the pronoun required in the task is the same as the case shown in the context. Examining model performance where a pronoun is shown in one case and then elicited in a different case would be an interesting area for further investigation, particularly given our findings that models have uncorrelated statistical biases across different cases.
More broadly, much of the recent progress on reasoning datasets has been critically investigated and shown to often be a result of spurious correlations and dataset artifacts (Trichelair et al., 2019). Although the results in this paper show models struggling with our task, we caution readers that our dataset actually gives a very generous estimate of model reasoning performance for two reasons. Firstly, accuracy is a generous metric compared to consistency-based metrics across parallel versions of each data instance that only vary the pronoun that we consider (Elazar et al., 2021a, b). As we randomly sample from our generated data, we do not have exactly parallel versions to compute this metric. Secondly, and more importantly, many of our task sentences are not “Google-proof” (Levesque et al., 2012), i.e., they can be solved with shallow heuristics such as word co-occurrences. Consider the following task sentence: The janitor said not to step on the wet floor, otherwise ___ would have to mop it all over again. The association between janitor and mop is strong compared to child (the participant paired with janitor) and mop, and could easily be exploited by models to solve the dataset without solving the task with something resembling “reasoning.”
Another shallow heuristic that can be used to solve our current dataset is to simply return the first pronoun in the context, which happens to always be the correct answer. Our dataset design is flexible and allows for the creation of other orderings of sentences, but this is another example of why our dataset should only be used as an evaluation dataset, and models should not be pre-trained or fine-tuned with any splits of our data, nor provided with examples for in-context learning.
We take steps to prevent data contamination following Jacovi et al. (2023), and do not evaluate with models behind closed APIs that do not guarantee that our data will not be used to train future models. However, as we cannot guarantee an absence of data leakage unless we never release the dataset, we encourage caution in interpreting results on our dataset with models trained on data after March 2024.
Acknowledgements
The authors thank Timm Dill for several rounds of patient annotation, Aaron Mueller for early feedback on this work, and Marius Mosbach and Vlad Niculae for discussions about math, plots and framing. Vagrant Gautam received funding from the BMBF’s (German Federal Ministry of Education and Research) SLIK project under the grant 01IS22015C. The work of Anne Lauscher is funded under the Excellence Strategy of the German Federal Government and States.
Appendix A List of occupations
The occupations along with their respective participants in parentheses are listed below in alphabetical order. This list is identical to the occupations and participants in Rudinger et al. (2018), except that we pair examiner with intern rather than victim:
accountant (taxpayer), administrator (undergraduate), advisor (advisee), appraiser (buyer), architect (student), auditor (taxpayer), baker (customer), bartender (customer), broker (client), carpenter (onlooker), cashier (customer), chef (guest), chemist (visitor), clerk (customer), counselor (patient), dietitian (client), dispatcher (bystander), doctor (patient), educator (student), electrician (homeowner), engineer (client), examiner (intern), firefighter (child), hairdresser (client), hygienist (patient), inspector (homeowner), instructor (student), investigator (witness), janitor (child), lawyer (witness), librarian (child), machinist (child), manager (customer), mechanic (customer) nurse (patient), nutritionist (patient), officer (protester), painter (customer), paralegal (client), paramedic (passenger), pathologist (victim), pharmacist (patient), physician (patient), planner (resident), plumber (homeowner), practitioner (patient), programmer (student), psychologist (patient), receptionist (visitor), salesperson (customer), scientist (undergraduate), secretary (visitor), specialist (patient), supervisor (employee), surgeon (child), teacher (student), technician (customer), therapist (teenager), veterinarian (owner), worker (pedestrian)
Appendix B Context Template Construction
For each case, we create 10 explicit templates which explicitly demonstrate the coreference between an individual and a pronoun using a subordinate clause, and 10 implicit templates, simple sentences which only contain a pronoun as the subject. The first two sentences of the context (i.e., the introduction and the first distractor) are always sampled from the explicit templates, and the rest are sampled from the implicit templates.
In both cases, we create five templates with terms with positive connotations (e.g., full, happy) and five opposing templates (i.e., hungry, unhappy). We denote $exp\_pos_{i}$ as the i-th positive explicit template where i ranges from 1 to 5, and $exp\_neg_{i}$ is the corresponding negative version. The introduction template can be selected from any of these 10 possibilities and filled with one of four pronouns.
After this, we pick a first distractor template, limiting ourselves to the five templates of the opposite sentiment of what we first picked, and also excluding the template of the same index and opposite polarity. For example, if we chose $exp\_pos_{3}$ as our introductory template, we would choose our first distractor template from $\{exp\_neg_{1},exp\_neg_{2},exp\_neg_{4},exp\_neg_{5}\}$.
After making a choice for the first distractor template, we fill it with any of the three remaining pronouns and then we remove this template’s index from our pool, but re-add the index of the introductory template. This is because subsequent distractor templates always use implicit templates. For example, if we chose $exp\_neg_{4}$ as our first distractor template, we would now choose from $\{imp\_neg_{1},imp\_neg_{2},imp\_neg_{5}\}$. For subsequent distractor templates, we sample from without replacement from these implicit templates.
Appendix C Annotation instructions
C.1 Task 1 Description
Together with this annotation protocol, you have received a link to a Google Sheet. The sheet contains 2 data columns and 2 task columns of randomized data. The data columns consist of
  • Sentences which you are asked to annotate for grammaticality; and
  • Questions about pronouns in the sentence, which you are asked to answer
Please be precise in your assignments and do not reorder the data. The columns have built-in data validation and we will perform further tests to check for consistent annotation.
C.1.1 Grammaticality
In the “Grammatical?” column, please enter your grammaticality judgments of the sentence, according to Standard English. The annotation options are:
  • grammatical (for fluent, syntactically valid and semantically plausible sentences)
  • ungrammatical (for sentences that have any typos, grammatical issues, or if the sentence describes a situation that don’t make sense, or just sounds weird)
  • not sure (if you are not sure whether it is clearly grammatical or ungrammatical)
Examples:
  • The driver told the passenger that he could pay for the ride with cash.
    => grammatical
  • The driver said the passenger that he could pay for the ride with cash.
    => ungrammatical (because ‘said’ is intransitive in Standard English)
C.1.2 Questions about pronouns
Every sentence contains a pronoun, and the “Question” column asks whether it refers to a person mentioned in the sentence or not. The annotation options are:
  • yes (if the pronoun refers to the person)
  • no (if the pronoun does not refer to the person)
  • not sure (if you are not sure about whether the pronoun refers to the person)
Examples:
  • The driver told the passenger that he could pay for the ride with cash.
    Does the pronoun he refer to the driver?
    => no
  • The driver told the passenger that he could pay for the ride with cash.
    Does the pronoun he refer to the passenger?
    => yes
C.2 Task 2 Description
Together with this annotation protocol, you have received a link to a Google Sheet. The sheet contains 1 randomized data column and 1 task column.
Each row in the data column consists of multiple sentences, of which precisely one sentence contains a blank. Your task is to determine the appropriate pronoun to fill in the blank, and enter it in the “Pronoun” column. Here, appropriate means correct in both form and case.
The tasks are designed to be unambiguous, so please provide only one solution and do not reorder the data.
Example:
  • The driver felt unhappy because he did not make enough money. The driver wondered whether ___ should take out a loan.
    => he
Appendix D Experimental Details
We use one 40GB NVIDIA A100 GPU for inference with most models, but we require two GPUs for OPT-30B and four for OPT-66B and the Llama-2-70B base and chat models. We access all models using the Huggingface Transformers library (Wolf et al., 2020) and use the minicons library for pseudo log likelihood evaluation (Misra, 2022).
Appendix E Prompting
Table 3 shows all 10 prompt templates we use with chat models. The first five are open-ended and the last five provide the model with pronoun options to choose from. Table 4 shows instantiated examples of one of the templates for both FLAN-T5 and Llama-2-chat. The number of maximum new tokens is set to 5 for FLAN-T5 and 20 for Llama-2-chat based on experiments.
ID Template
[…]
Table 3: Prompting templates for chat models, where “task” is filled with each dataset instance, and “options” is a list of four pronouns to choose from, all in the correct case.
Model type Example input
FLAN-T5 Fill in the blank with the appropriate pronoun.
The technician told the customer that ___ could pay with cash.
OPTIONS:
- he
- she
- they
- xe
Llama-2-chat [INST] Fill in the blank with the appropriate pronoun.
The technician told the customer that ___ could pay with cash.
OPTIONS:
- he
- she
- they
- xe [/INST]
Table 4: Example input using template 8 for FLAN-T5 and Llama-2-chat models.
Appendix F Model correlations between pronoun predictions by case
Table 5 shows correlations between pronoun predictions across cases for individual models. The distribution of these scores are visualized in Figure 2.
Model Nom-Acc Nom-Poss Acc-Poss
BERT-base 0.32 0.21 0.31
BERT-large -0.05 0.35 0.03
RoBERTa-base -0.46 0.24 -0.27
RoBERTa-large 0.26 0.37 0.23
ALBERT-base-v2 0.19 -0.16 0.27
ALBERT-large-v2 0.06 0.07 0.22
ALBERT-xlarge-v2 0.26 0.30 0.03
ALBERT-xxlarge-v2 0.41 0.28 0.18
MosaicBERT -0.23 0.21 -0.32
OPT-125M -0.03 0.42 0.03
OPT-350M -0.04 0.35 0.04
OPT-1.3B -0.25 0.39 -0.18
OPT-2.7B -0.12 0.15 -0.27
OPT-6.7B 0.00 0.20 -0.10
OPT-13B -0.08 0.48 -0.07
OPT-30B -0.32 0.23 -0.01
OPT-66B 0.17 0.29 0.17
Pythia-14M 0.12 0.23 0.10
Pythia-70M 0.32 0.29 0.05
Pythia-160M 0.47 0.40 0.03
Pythia-410M 0.41 0.30 0.05
Pythia-1B -0.12 0.49 -0.14
Pythia-1.4B 0.15 0.35 0.11
Pythia-2.8B -0.16 0.42 -0.14
Pythia-6.9B -0.12 0.53 0.06
Pythia-12B 0.02 0.54 -0.07
Llama-2-7B 0.09 0.34 -0.21
llama-2-13B 0.00 0.55 -0.01
llama-2-70B 0.36 0.47 0.02
Table 5: Individual correlations between pronoun predictions across different cases.