### v2

# Robust Pronoun Fidelity with English LLMs: Are they Reasoning, Repeating, or Just Biased?

Abstract
Robust, faithful and harm-free pronoun use for individuals is an important goal for language models as their use increases, but prior work tends to study only one or two of these characteristics at a time. To measure progress towards the combined goal, we introduce the task of pronoun fidelity: given a context introducing a co-referring entity and pronoun, the task is to reuse the correct pronoun later. We present RUFF, a carefully-designed dataset of over 5 million instances to measure robust pronoun fidelity in English, and we evaluate 37 popular large language models across architectures (encoder-only, decoder-only and encoder-decoder) and scales (11M-70B parameters). When an individual is introduced with a pronoun, models can mostly faithfully reuse this pronoun in the next sentence, but they are significantly worse with she/her/her, singular they and neopronouns. Moreover, models are easily distracted by non-adversarial sentences discussing other people; even one additional sentence with a distractor pronoun causes accuracy to drop on average by 34%. Our results show that pronoun fidelity is neither robust, nor due to reasoning, in a simple, naturalistic setting where humans achieve nearly 100% accuracy. We encourage researchers to bridge the gaps we find and to carefully evaluate reasoning in settings where superficial repetition might inflate perceptions of model performance.
1 Introduction
Third-person pronouns (he, she, they, etc.) are words that construct individuals’ identities in conversations (Silverstein, 1985). In English, these pronouns mark referential gender for the entity they are referring to, which can also index an individual’s social gender, e.g., man, woman, non-binary (Cao and Daumé III, 2020). Correctly using the pronouns an individual identifies with is important, as misgendering (including through incorrect pronoun use) can in the best case be a social faux pas (Stryker, 2017) and in the worst case, cause psychological distress, particularly to transgender individuals (McLemore, 2018).
Accordingly, it is important for large language models (LLMs) to use pronouns faithfully and without causing harm. To this end, many studies have explored how LLMs handle pronouns, showing that they stereotypically associate pronouns and occupations (Kurita et al., 2019), reason about co-referring pronouns and entities better when they conform to stereotypes (Tal et al., 2022), fail when exposed to novel pronoun phenomena such as neopronouns (Lauscher et al., 2023), and cannot consistently reuse neopronouns during generation (Ovalle et al., 2023a). These shortcomings create quality of service differences and cause representational harm, amplifying discrimination against certain pronoun users (Blodgett et al., 2020; Dev et al., 2021).
However, a question that has gone unexamined thus far is: How robust is model faithfulness to pronouns when discussing more than one person? To answer this question, we propose pronoun fidelity2), a new task to investigate realistic model reasoning about pronouns, and we introduce RUFF3), a novel, large-scale dataset of over 5 million instances, carefully designed to evaluate this task. With this dataset, we present an analysis of pronoun fidelity across 37 popular language models covering a range of architectures and scales (11M-70B), to investigate whether models are reasoning, repeating, or just biased.
First, we collect model pronoun predictions for occupations in the absence of context, to establish a “bias baseline” (§5). Next, we evaluate whether models can overcome their biased pronoun predictions when explicitly shown what pronoun to use in context (§6). All models are good at this task, but there are significant disparities across pronoun sets. We then test the robustness of this result by inserting naturalistic distractor sentences using a different pronoun to talk about another person (§7). Even one non-adversarial distractor sentence vastly deteriorates model performance as shown in Figure 1. Finally, in a detailed error analysis (§8), we disentangle whether model errors can be attributed to distraction or falling back to bias, finding that encoder-only and decoder-only models behave in fundamentally different ways.
Figure 1: We evaluate model accuracy at using the correct pronoun for an entity when provided with an explicit introduction and 0-5 non-adversarial distractor sentences. Llama-2-70B and RoBERTa-large show large accuracy drops with just one distractor. Accuracy is averaged over 3 data splits; standard deviation is shown with shading.
Overall, our results show that models struggle to reason about pronouns in a simple, naturalistic setting and highlight the need for careful task design to ensure that superficial repetition does not lead to inflated claims about model reasoning. We release all code and data to encourage researchers to bridge the gaps we find: https://github.com/uds-lsv/robust-pronoun-fidelity
2 Pronoun Fidelity Task
Discussing multiple individuals is natural, frequent and well-studied in discourse; we use both definite references and pronouns in natural language to establish continuity and coherence (Grosz et al., 1995). We formalize a version of these phenomena in our task: given a context in which a co-referring entity and pronoun are introduced, the task is to reconstruct the pronoun later in a sentence about the entity, independent of a limited number of potential distractors.
[…]
Introduction: The accountant had just eaten a big meal so her stomach was full.
Distractor 1: The taxpayer needed coffee because their day had started very early.
Distractor N: Their sleep had been fitful.
Task sentence: The accountant was asked about ___ charges for preparing tax returns.
[…]
the probability $P$ of reconstructing the correct pronoun $p_{a}$ in the sentence $t({\color[rgb]{0.13,0.55,0.13}\definecolor[named]{pgfstrokecolor}{rgb}{0.13,0.55,0.13}e_{a}},p)$, given the context.
3 RUFF Dataset
RUFF, our evaluation dataset for Robust pronoUn Fidelity at scale, is constructed with a 3-step pipeline to create naturalistic instances of our task with unambiguous answers: template creation (§3.1), assembly (§3.2), and validation (§3.3).
Our data covers 60 occupations (see Appendix B) and four third-person pronouns (he, she, they and xe) in three grammatical cases (nominative, accusative and possessive dependent). Our occupations (and corresponding participants for distractor sentences) are chosen from Winogender schemas (Rudinger et al., 2018), as their bias characteristics are well-studied in NLP. In addition to the English masculine (he/him/his) and feminine (she/her/her) pronouns, we heed Lauscher et al.’s (2022) call for more inclusive NLP research by examining two more pronoun sets that are less well-studied in NLP: singular they (they/them/their), the pronoun of choice of over 75% of respondents to the Gender Census (Lodge, 2023), and xe/xem/xyr, the most popular neopronoun according to the same census.
3.1 Template creation
Task templates. We create one task sentence template per occupation and grammatical case for a total of 180 templates, designed to create an unambiguous coreference with the occupation only. For instance, charges for preparing tax returns can only belong to an accountant, never a taxpayer, which is the corresponding participant.
Context templates. We create explicit (definite reference + pronoun) and implicit (pronoun-only) versions of 10 context templates per grammatical case, for a total of 30 templates. Each explicit context template begins with an entity and introduces the pronoun in a clause, e.g., The taxpayer needed coffee because their day had started very early, while implicit templates are simple sentences like Their sleep had been fitful. All templates contain generic themes, e.g., universal human emotions and sensations (hungry/full, tired/energetic, unhappy/happy, etc.). These templates can thus be applied to all occupations and participants, and allow for a controlled but still coherent setting.
3.2 Template assembly
We then instantiate and combine the previously created templates to assemble our data instances. First, we select an occupation ($e_{a}$) and one of its task templates. We pick a pronoun ($p_{a}$) to use as ground truth and instantiate a random context template with the selected occupation and pronoun. The simplest version of the pronoun fidelity task includes just this introduction sentence followed by the task sentence. Instantiating 10 templates with 4 different pronoun sets and pairing them with task templates for 60 occupations across 3 grammatical cases gives us a total of 7,200 unique instances for this version of the task.
To create more complex data instances, we insert a variable number of distractor sentences between the introduction and task sentences, discussing a participant $e_{b}$ who uses a different pronoun $p_{b}$. These are also sampled from the set of context templates (see Appendix C for details). Instantiating 4 templates with 3 unused pronouns gives 86,400 unique instances with one distractor.
Data typespace spaceNumber of instances
With no context
Task sentences 180
With introductory context
+ 0 distractors 3 x 2,160 0,00(of 7,200)
+ 1 distractors 3 x 2,160 0,0(of 86,400)
+ 2 distractors 3 x 2,160 0,(of 345,600)
+ 3 distractors 3 x 2,160 (of 1,036,800)
+ 4 distractors 3 x 2,160 (of 2,073,600)
+ 5 distractors 3 x 2,160 (of 2,073,600)
Table 1: Number of dataset instances. Pronoun fidelity instances consist of task instances combined with introductory contexts and optional distractors. We subsample 3 sets of 2,160 sentences (of the total number of instances we created).
Our stackable dataset design allows us to generate a vast amount of data of varying lengths, giving us a controlled setting to evaluate context effects on model predictions. We subsample the data with three random seeds for the rest of our evaluation, ensuring that all occupations, cases, pronoun declensions and distractor pronouns are equally represented in each subsampled set of 2,160 sentences. All data statistics are shown in Table 1.
Model Sizes Architecture
Evaluated with (Pseudo) Log Likelihoods
ALBERT-v2 base (11M), large (17M), xlarge (58M), xxlarge (223M) Encoder-only
BERT base (110M), large (340M) Encoder-only
RoBERTa base (125M), large (355M) Encoder-only
MosaicBERT 137M Encoder-only
OPT 125M, 350M, 1.3B, 2.7B, 6.7B, 13B, 30B, 66B Decoder-only
Pythia 14M, 70M, 160M, 410M, 1B, 1.4B, 2.8B, 6.9B, 12B Decoder-only
Llama-2 7B, 13B, 70B Decoder-only
Evaluated with prompting
FLAN-T5 small (77M), base (248M), large (783), xl (2.85B), xxl (11.3B) Encoder-decoder
Llama-2-chat 7B, 13B, 70B Decoder-only
Table 2: Models we experiment with across a range of sizes and architectures.
3.3 Data validation
We validate all task and context templates. To verify that the pronoun fidelity task is easy and unambiguous for humans, we also validate a subset of these instances with 0-5 distractors. Annotator information is shown in Appendix D and all annotator instructions are provided in Appendix E.
Templates. Two authors with linguistic training iteratively created and validated sentence templates for grammaticality and correct coreferences until consensus was reached. An additional annotator independently rated 100% of the sentences as grammatical and with the correct coreferences.
Pronoun fidelity task. We sampled 100 instances with each possible number of distractors (0-5), for a total of 600 instances. One author and one annotator had to fill in the pronoun and they each performed with 99.8% accuracy. [^1] They disagreed on non-overlapping instances which appeared to be random slips.
4 Experimental Setup
We list our models, evaluation methods, and metrics. Further details are provided in Appendix F.
4.1 Models
We experiment with 37 transformer-based language models (see Table 2), chosen to evaluate the effects of architecture and scaling. Our 9 encoder-only models are from the BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), ALBERT-v2 (Lan et al., 2020) and MosaicBERT (Portes et al., 2023) model families, as the first three remain well-used in NLP, and the last is trained on much more data. As for our 20 decoder-only models, we select the popular Llama-2 (Touvron et al., 2023) model family, as well as OPT (Zhang et al., 2022) and Pythia (Biderman et al., 2023) for their large range of model sizes. In Appendix I, we also experiment with eight popular chat models that are further trained with instruction-tuning and reinforcement learning, to evaluate task performance with prompting; specifically, we use decoder-only Llama-2-chat models (Touvron et al., 2023) and encoder-decoder FLAN-T5 models (Chung et al., 2022).
4.2 Obtaining predictions
Decoder-only models. For the majority of our experiments, we follow Hu and Levy (2023) in taking direct measurements of probabilities as a proxy for models’ metalinguistic judgements. We verbalize four versions of each data instance, i.e., we fill in the blank with each of the four pronouns we consider, creating four options. We then compute instance-level model log likelihoods for these options, and select the option with the highest log likelihood as the model’s choice.
Encoder-only models. For comparability with decoder-only models, we follow the same procedure for encoder-only models, but use pseudo log likelihoods (Salazar et al., 2020; Kauf and Ivanova, 2023). We do not use masked token prediction due to tokenization issues with neopronouns (Ovalle et al., 2023b); briefly, we want xe to be tokenized “normally” (which is often as two tokens) rather than a single UNK token.
Chat models. Following common practice, we evaluate chat models (FLAN-T5 and Llama-2-chat) using vanilla and chain-of-thought prompting in Appendix I. Following Sclar et al. (2024), we show the range of expected performance with 10 different prompts (see Appendix G), inspired by the prompts to elicit coreferences in the FLAN collection (Longpre et al., 2023).
4.3 Metrics
As every instance of the pronoun fidelity task has a unique correct answer, we report accuracy averaged over the three randomly sampled subsets of our dataset. We show the standard deviation with error bars or shading. Where possible, we perform significance testing with a Welch’s t-test and a threshold of 0.05. We use human performance as our ceiling, and compare models to a baseline of randomly selecting 1 of the 4 pronouns (i.e., 25%).
Figure 2: Counts of pronoun predictions from all models, in the absence of context. Error bars indicate standard deviation across models.
5 Model Predictions with No Context
We begin by creating a “bias baseline,” i.e., obtaining pronoun predictions from models on our task sentences in the absence of any context. In Section 6, we will examine whether models can overcome this bias with reasoning when provided with context establishing a single correct answer.
[…]
Example: The accountant was asked about ___ charges for preparing tax returns.
[…]
As we cannot evaluate accuracy on a task with no single correct answer, we show the counts of model predictions of different pronoun declensions in Figure 2, averaged over all models. Model-specific counts are shown in Appendix H. Even though our task sentences are designed such that any pronoun set can be used grammatically, all models tend to assign higher probability to he/him/his than other pronoun sets.
Obtaining pronoun predictions without context is a popular method to measure model bias, with numerous papers (Kurita et al., 2019, inter alia) showing that associations between occupations and pronouns are based on social gender stereotypes, e.g., doctor-he and nurse-she. However, model pronoun predictions might reflect dataset artifacts such as the choice of occupations, or be a statistical accident of the chosen templates (Seshadri et al., 2022). In addition, intrinsic biases may not correlate with actual pronoun use with context (Goldfarb-Tarrant et al., 2021). In order to test for such extrinsic behaviours, the rest of this paper examines whether models can override their intrinsic statistical biases on these same templates when provided with the right pronoun to use.
Figure 3: Pronoun fidelity by model with an introductory context. Accuracy is averaged across occupations, pronouns and grammatical cases, and is above chance (0.25) but below human performance (1.0).
he/him/his Refer to caption
she/her/her
(* < he/him/his)
they/them/their
(* < she/her/her)
xe/xem/xyr
(* < she/her/her)
Figure 4: Pronoun fidelity by model with an introductory context, split by pronoun series. Model accuracy is compared to chance (0.25) and human performance (1.0) and * denotes statistical significance.
6 Injecting an Introductory Context
When models are provided with an introductory sentence explicitly establishing the pronoun to use for an entity, can they use that pronoun to refer to the same entity in the immediate next sentence?
[…]
Example: The accountant had just eaten a big meal so her stomach was full. The accountant was asked about ___ charges for preparing tax returns.
[…]
As Figure 3 shows, all models perform better than chance at pronoun fidelity with a simple introduction (up to 0.95 with MosaicBERT), but not as well as humans, who achieve perfect performance. We also see improvements with increasing model scale, with the exception of ALBERT-v2, as in Tay et al. (2022).
Which pronouns are harder? Even in the simplest case of the pronoun fidelity task, patterns emerge when split by pronoun, as shown in Figure 4. Overall model accuracy on he/him/his is significantly higher than she/her/her, which in turn is significantly higher than both they/them/their and xe/xem/xyr, in line with previous findings that language technology has gaps when it comes to neopronouns (Lauscher et al., 2023). Models show intriguing patterns with these last two pronoun sets. Most encoder-only models appear to handle the neopronoun better than singular they (e.g., BERT-large has an accuracy of 0.78 on xe/xem/xyr compared to 0.60 on they/them/their), which warrants further investigation. Decoder-only models smaller than 6.7B parameters struggle with the neopronoun, with every OPT and Pythia model smaller than 2.7B parameters performing below chance, and in some cases (e.g., Pythia-14M, Pythia-70M and Pythia-160M) even performing close to 0.0. Beyond this scale, however, models perform better on xe/xem/xyr than on singular they, with Llama-13B achieving 0.96 accuracy on the neopronoun. These differences are statistically significant. As the training data for individual model families is the same, this might suggest that decoder-only models generalize to novel pronouns starting at the scale of 6.7B parameters, but in light of Schaeffer et al. (2023), this result could just as well be a mirage resulting from our use of accuracy, a discontinuous metric. In either case, our observations could also explain the poor performance that some previous studies of neopronouns find, as the largest model that Hossain et al. (2023) experiment with, for instance, is OPT-6.7B. The lower performance of bigger models with singular they could also be a reflection of human processing difficulties with definite, specific (singular) they, as has been observed in linguistics (Conrod, 2019).
(a) OPT
(b) Pythia
(c) Llama-2
(d) ALBERT-v2
(e) BERT
(f) RoBERTa
Figure 5: With more distractors, decoder-only models (above) get steadily worse; encoder-only models (below) get worse with one distractor and then recover, plateauing below their no-distractor accuracy.
(a) With 0 distractors
(b) With 5 distractors
Figure 6: Scaling behaviour by architecture. With 0 distractors (above), encoder-only models are comparable to decoder-only models orders of magnitude larger. With 5 distractors (below), encoder-only models are far better.
7 Adding Distractors
To further probe whether models actually “reason” when provided with context, we systematically inject sentences containing distractor pronouns between the introduction and the task, reflecting a natural usage scenario where multiple people are discussed with definite references and pronouns.
[…]
Example: The accountant had just eaten a big meal so her stomach was full. The taxpayer needed coffee because their day had started very early. Their sleep had been fitful. The accountant was asked about ___ charges for preparing tax returns.
Correct answer: her
[…]
Figure 5 shows that distractors degrade performance for all models. Encoder-only and decoder-only models show different performance curves as more distractors are added: all decoder-only models get steadily worse, whereas encoder-only models perform the worst with one distractor and then seem to slowly recover, never quite reaching their level of performance with no distractors. Scaling generally holds within model families, with larger models performing better with more distractors than smaller models of the same type. Figure 6 examines the interplay of scaling and architecture at a higher level, comparing results on the easiest case of pronoun fidelity (no distractors) with the hardest case (5 distractors). Surprisingly, with no distractors, encoder-only models are much better than decoder-only models of the same scale, and their performance is comparable to or better than decoder-only models that are orders of magnitude larger; RoBERTa-base (125M) is 0.86 accurate compared to OPT-125M’s 0.55, and exceeds OPT-66B’s 0.83 despite being more than 500 times smaller. In the hardest version of our task with five distractors, encoder-only models are far better than all decoder-only models, which show dramatically degraded performance; Llama-70B only achieves 0.37 accuracy, compared to MosaicBERT’s impressive 0.87. The lack of robustness of decoder-only models to distractors is striking, given that most state-of-the-art models today are decoder-only models. We hypothesize that architectural differences might explain the performance gaps; encoder-only models might use bidirectional attention to more closely relate the entity mentions in the introduction and task sentences. Training on next token prediction might also make decoder-only models prone to recency bias.
Using vanilla and chain-of-thought prompting (Appendix I) show the same patterns of degradation, reinforcing that model pronoun fidelity is not robust, and good performance with no distractors (§6) is likely not due to “reasoning” at all.
(a) OPT
(b) Pythia
(c) Llama-2
(d) ALBERT-v2
(e) BERT
(f) RoBERTa
Figure 7: Trends in model distractibility (use of the distractor pronoun) and model bias (reverting to the context-free prediction). With more distractors, the proportion of errors due to distraction increases for decoder-only models (above) and decreases for encoder-only models (below).
8 Distractibility versus Bias
In adding distractor sentences, we add distance from the introduction via additional tokens that might make the model forget the original occupation-pronoun association, and the distractor pronoun also acts as a competing token that the model might accidentally repeat. In this section, we focus on the error cases to disentangle whether models are “forgetting” and reverting to biased predictions from Section 5, or if they are actually being distracted. When a model gets the answer wrong, it is for one of three reasons: (1) distractibility, i.e., repeating the distractor pronoun, (2) bias, i.e., reverting to the model’s context-free prediction, or (3) picking a completely different pronoun. All three possibilities are illustrated below , and we hypothesize that the first two possibilities are much more frequent than the third.
In cases where the distractor pronoun is the same as the model’s context-free prediction, it is impossible to disentangle distractibility and bias just from the model’s prediction. Hence, we exclude these and focus on the unambiguous error cases. As expected, we find that 74-93% of unambiguous model errors can be attributed to either model distractibility or bias.
[…]
Context-free5)
Example: The accountant was asked about ___ charges for preparing tax returns.
Prediction: his
With introduction and distractors7)
Example: The accountant had just eaten a big meal so her stomach was full. The taxpayer needed coffee because their day had started very early. Their sleep had been fitful. The accountant was asked about ___ charges for preparing tax returns.
Correct answer: her
Distraction error: their
Bias error: his
Other error: xyr
[…]
We first examine model distractibility, i.e., what percentage of errors are caused by models repeating the distractor pronoun instead of the correct pronoun. As expected, Figure 7 shows that across models, distraction is indeed the primary type of error for most models. Decoder-only models get increasingly distractible with more distractors, i.e., the proportion of errors due to distractor pronoun repetition steadily increases as distractors are added, saturating just below 85%. On the other hand, encoder-only models seem to become less distractible with the addition of more distractors. We know from the previous section that encoder-only models recover in their pronoun fidelity with 2-5 distractors, but here we measure distractibility as a percentage of all errors. Thus, a constant or increasing proportion of all the model errors could be due to distraction, and the fact that it isn’t for encoder-only models is quite surprising! We leave it to future work to investigate whether this behaviour relates to positional bias or context use.
As their proportion of distraction errors goes down, encoder-only models increasingly revert to biased predictions. With BERT-large in particular, as soon as there is more than one distractor, the biggest proportion of errors is due to bias rather than distraction. BERT-large appears more biased and less distractible than BERT-base, in contrast to all other models. Generally, larger models seem to be more distractible and revert to their bias less often, whereas smaller models are more biased and less distractible. Our findings on bias errors contrast with Tal et al. (2022), where larger models make a higher proportion of bias errors on a downstream task than smaller models. This might be due to our task having distractors, which seem to strongly influence model behaviour in this setting.
The high distractibility of all models strongly suggests that models simply repeat the most recent pronoun, rather than using anything resembling “reasoning,” and the contrast in error behaviour between encoder-only and decoder-only models further highlights their differences. This shows that claims about decoder-only models should not be applied to all LLMs, and that reasoning must be evaluated carefully, accounting for the possibility of inflated performance due to shallow heuristics like repetition.
9 Discussion and Future Work
Our results show that even the biggest models of today are not up to the task of pronoun fidelity once it includes a single sentence discussing another person. All models are easily distracted, but encoder-only models and decoder-only models show very different patterns both in performance degradation with more distractors and their reasons for errors. Performance on this type of reasoning task should be evaluated carefully, with attention to how the overall patterns break down by different pronouns and accounting for the possibility of repetition. Below we expand on some questions raised by our findings.
Improving robust pronoun fidelity. A natural direction of future work is to solve the problem of robust pronoun fidelity. We urge researchers interested in this direction to treat RUFF as an evaluation dataset, as it was designed. Due to the presence of positional and associative heuristics that we expand on in Appendix A, RUFF should not be seen as a source of data for fine-tuning or in-context learning, which is also why we do not run these experiments. Instead, a promising direction might be to encourage models to explicitly track associations between pronoun sets and individuals, just as people do.
On “reasoning.” Throughout the paper, we refer to “reasoning,” but this is inaccurate. What looks like “reasoning” when we inject an introductory context starts to look much more like repetition—or stochastic parroting (Bender et al., 2021)—when we add a distractor and see the same models performing drastically worse. Even the higher performance of encoder-only models cannot accurately be attributed to “reasoning” in the same way that we use this word for humans, as these models are not grounded in meaning from the real world (Bender and Koller, 2020). We use the word reasoning in line with other work in the field, but note that as these are all language models, it is more accurate to say that the way that decoder-only models model language is prone to repetition of recent examples of the same word class, compared to encoder-only models.
Why exactly do we see the patterns we see? Our dataset design and error analysis shed light on model behaviour, allowing us to evaluate different architectures comparably and disentangle the effects of repetition, distraction and statistical bias. However, it is beyond the scope of this paper to investigate where in the model architecture, neurons or pre-training data this comes from and what we can do about it towards improving reasoning and mitigating bias. Tools from interpretability literature, e.g., attribution analysis, could help here, and are an important direction for future work.
Beyond our dataset. Given the breadth of our task definition, future work could include examining pronoun fidelity for participants, for names by extending Hossain et al. (2023), with differently ordered sentences, and in more natural settings such as stories (Mostafazadeh et al., 2016) or with real-world data as in Webster et al. (2018) and Levy et al. (2021). Additionally, we evaluate on a version of this task that allows us to quantify repetition, i.e., the grammatical case of the elicited pronoun is the same as the case shown in the context. Examining model performance where a pronoun is shown in one grammatical case and then elicited in a different one would be interesting to probe syntactic generalization.
10 Related Work
Pronoun fidelity.Hossain et al. (2023) and Ovalle et al. (2023a) both study pronoun fidelity when models are prompted with a pronoun series to use for an individual, but they only consider simplistic pronoun use with no more than one person at a time. Although we look at within-language pronoun use, faithful pronoun use in context has also been studied in machine translation (Müller et al., 2018; Voita et al., 2018; Fernandes et al., 2023), where there is also a ground truth. Similar to our work, Sharma et al. (2022) injects context with an explicit coreference to encourage faithful pronoun translation. However, none of these papers explore the robustness of pronoun fidelity in the presence of distractors.
Reasoning with pronouns. Most existing work about LLM reasoning with pronouns focuses on the task of coreference resolution, i.e., the ability to identify the connection between a pronoun and an entity, which may not translate to faithful reuse of that pronoun later, as in our work. Reasoning with pronouns typically uses Winograd schemas (Levesque et al., 2012; Abdou et al., 2020; Emelin and Sennrich, 2021), or Winograd-like schemas about named individuals (Webster et al., 2018; Zhao et al., 2018), or people referred to by their occupation (Rudinger et al., 2018; Levy et al., 2021). Most studies focus on he and she, but recent work has expanded to include singular they (Baumler and Rudinger, 2022) and with neopronouns (Cao and Daumé III, 2021; Felkner et al., 2023), as we do.
Pronouns and occupational bias. Stereotypical associations between pronouns and occupations have been studied in masked token prediction (Kurita et al., 2019; de Vassimon Manela et al., 2021; Tal et al., 2022) and embeddings (Bolukbasi et al., 2016; Zhao et al., 2019), but these studies typically use brittle methodology (Gonen and Goldberg, 2019; Seshadri et al., 2022) and measure intrinsic bias, which may not translate to extrinsic bias or harms (Goldfarb-Tarrant et al., 2021). Unlike these works, we evaluate extrinsic bias and performance through our focus on natural pronoun use in context.
Robustness in context. The impact of context on the robustness of language model reasoning has been investigated in many areas other than pronoun fidelity, e.g., negation (Gubelmann and Handschuh, 2022), linguistic acceptability (Sinha et al., 2023), natural language inference (Srikanth and Rudinger, 2022), and question answering  (Liu et al., 2024; Levy et al., 2024).
11 Conclusion
We introduce the task of pronoun fidelity to evaluate robust, faithful and harm-free pronoun use in language models, and we present RUFF, a dataset we designed to evaluate it. We find evidence of faithful pronoun use only in a very simple setting, i.e., where only one person is discussed. Even here, models show significant performance disparities with neopronouns, singular they and she/her/her, compared to he/him/his. Even adding a single sentence about a second individual with a different pronoun causes accuracy to drop dramatically, showing that pronoun fidelity is neither robust to non-adversarial distractors nor due to “reasoning.” As more distractor sentences are added, encoder-only models perform better overall, but increasingly revert to biased predictions, while decoder-only models get increasingly distracted. Our results show that in a setting that is very simple for humans, widely-used large language models are unable to robustly and faithfully reason about pronouns, and continue to amplify discrimination against users of certain pronouns. We encourage researchers to bridge the performance gaps we report and to more carefully evaluate “reasoning,” especially when simple repetition could inflate perceptions of model performance.
Acknowledgements
The authors thank Timm Dill for several rounds of patient annotation, as well as Aaron Mueller, Marius Mosbach, Vlad Niculae and Yanai Elazar, for feedback and discussions about math, plots and framing. Vagrant Gautam received funding from the BMBF’s (German Federal Ministry of Education and Research) SLIK project under the grant 01IS22015C. The work of Anne Lauscher is funded under the Excellence Strategy of the German Federal Government and States.
Appendix A Limitations
Shallow heuristics. Much of the recent progress on reasoning datasets has been critically investigated and shown to often be a result of spurious correlations and dataset artifacts (Trichelair et al., 2019; Elazar et al., 2021). We caution readers that our dataset also gives a very generous estimate of model reasoning performance for multiple reasons. Many of our task sentences are not “Google-proof” (Levesque et al., 2012), i.e., they can be solved with shallow heuristics such as word co-occurrences. Consider the following task sentence: The janitor said not to step on the wet floor, otherwise ___ would have to mop it all over again. The association between janitor and mop is strong compared to child (the participant paired with janitor) and mop, and could easily be exploited by models to solve the dataset without solving the task with something resembling “reasoning.” Another shallow heuristic that can be used to solve our current dataset is to simply return the first pronoun in the context, which happens to always be the correct answer. Our dataset design is flexible and allows for the creation of other orderings of sentences, but this is another example of why our dataset in its current form should only be used as an evaluation dataset, and models should not be pre-trained or fine-tuned with any splits of our data, nor provided with examples for in-context learning.
Whose bias? Our task as it is defined in Section 2 is much broader than the scope of our dataset. We focus on occupations due to the wide attention they have received in prior literature, but we continue a long tradition of ignoring biases relating to the participants, e.g., child, taxpayer, etc.
Data contamination. We take steps to prevent data contamination following Jacovi et al. (2023), including not releasing our data in plain text, and not evaluating with models behind closed APIs that do not guarantee that our data will not be used to train future models. However, as we cannot guarantee a complete absence of data leakage unless we never release the dataset, we encourage caution in interpreting results on RUFF with models trained on data after March 2024.
Appendix B List of Occupations
The occupations along with their respective participants in parentheses are listed below in alphabetical order. This list is identical to the occupations and participants in Rudinger et al. (2018), except that we pair examiner with intern rather than victim:
accountant (taxpayer), administrator (undergraduate), advisor (advisee), appraiser (buyer), architect (student), auditor (taxpayer), baker (customer), bartender (customer), broker (client), carpenter (onlooker), cashier (customer), chef (guest), chemist (visitor), clerk (customer), counselor (patient), dietitian (client), dispatcher (bystander), doctor (patient), educator (student), electrician (homeowner), engineer (client), examiner (intern), firefighter (child), hairdresser (client), hygienist (patient), inspector (homeowner), instructor (student), investigator (witness), janitor (child), lawyer (witness), librarian (child), machinist (child), manager (customer), mechanic (customer) nurse (patient), nutritionist (patient), officer (protester), painter (customer), paralegal (client), paramedic (passenger), pathologist (victim), pharmacist (patient), physician (patient), planner (resident), plumber (homeowner), practitioner (patient), programmer (student), psychologist (patient), receptionist (visitor), salesperson (customer), scientist (undergraduate), secretary (visitor), specialist (patient), supervisor (employee), surgeon (child), teacher (student), technician (customer), therapist (teenager), veterinarian (owner), worker (pedestrian)
Appendix C Context Template Construction
Semantically bleached templates such as He is an accountant are well-established in the literature for testing word embedding associations (Caliskan et al., 2017), but it is unnatural to use more than one consecutively, which is necessary to evaluate robustness in this setting. On the other hand, natural corpora like Levy et al. (2021) contain occupation-specific sentences that cannot be used more generally for a controlled setting across occupations and participants. As a compromise between a controlled setting and naturalistic data, we create context templates with generic themes that apply to all humans, e.g., universal emotions and sensations (hungry/full, tired/energetic, unhappy/happy, etc.).
For each case, we create 10 explicit templates which explicitly demonstrate the coreference between an individual and a pronoun using a subordinate clause, and 10 implicit templates, simple sentences which only contain a pronoun as the subject. The first two sentences of the context (i.e., the introduction and the first distractor) are always sampled from the explicit templates, and the rest are sampled from the implicit templates, as this reflects natural and coherent use of pronouns in discourse.
In both cases, we create five templates with terms with positive connotations (e.g., full, happy) and five opposing templates (i.e., hungry, unhappy). We denote $exp\_pos_{i}$ as the i-th positive explicit template where i ranges from 1 to 5, and $exp\_neg_{i}$ is the corresponding negative version. The introduction template can be selected from any of these 10 possibilities and filled with one of four pronouns.
After this, we pick a first distractor template, limiting ourselves to the five templates of the opposite sentiment of what we first picked, and also excluding the template of the same index and opposite polarity. For example, if we chose $exp\_pos_{3}$ as our introductory template, we would choose our first distractor template from $\{exp\_neg_{1},exp\_neg_{2},exp\_neg_{4},exp\_neg_{5}\}$.
After making a choice for the first distractor template, we fill it with any of the three remaining pronouns and then we remove this template’s index from our pool, but re-add the index of the introductory template. This is because subsequent distractor templates always use implicit templates. For example, if we chose $exp\_neg_{4}$ as our first distractor template, we would now choose from $\{imp\_neg_{1},imp\_neg_{2},imp\_neg_{5}\}$. For subsequent distractor templates, we sample without replacement from these implicit templates.
Appendix D Annotator Demographics
All three annotators (two authors and an additional annotator) are fluent English speakers. The two authors who create and validate templates have linguistic training at the undergraduate level. One author and one annotator have experience with using singular they and neopronouns, while the other author has prior exposure to singular they but not the neopronoun xe.
Appendix E Annotation Instructions
E.1 Task 1 Description
Together with this annotation protocol, you have received a link to a Google Sheet. The sheet contains 2 data columns and 2 task columns of randomized data. The data columns consist of
  • Sentences which you are asked to annotate for grammaticality; and
  • Questions about pronouns in the sentence, which you are asked to answer
Please be precise in your assignments and do not reorder the data. The columns have built-in data validation and we will perform further tests to check for consistent annotation.
E.1.1 Grammaticality
In the “Grammatical?” column, please enter your grammaticality judgments of the sentence, according to Standard English. The annotation options are:
  • grammatical (for fluent, syntactically valid and semantically plausible sentences)
  • ungrammatical (for sentences that have any typos, grammatical issues, or if the sentence describes a situation that don’t make sense, or just sounds weird)
  • not sure (if you are not sure whether it is clearly grammatical or ungrammatical)
Examples:
  • The driver told the passenger that he could pay for the ride with cash.
    => grammatical
  • The driver said the passenger that he could pay for the ride with cash.
    => ungrammatical (because ‘said’ is intransitive in Standard English)
E.1.2 Questions about pronouns
Every sentence contains a pronoun, and the “Question” column asks whether it refers to a person mentioned in the sentence or not. The annotation options are:
  • yes (if the pronoun refers to the person)
  • no (if the pronoun does not refer to the person)
  • not sure (if you are not sure about whether the pronoun refers to the person)
Examples:
  • The driver told the passenger that he could pay for the ride with cash.
    Does the pronoun he refer to the driver?
    => no
  • The driver told the passenger that he could pay for the ride with cash.
    Does the pronoun he refer to the passenger?
    => yes
E.2 Task 2 Description
Together with this annotation protocol, you have received a link to a Google Sheet. The sheet contains 1 randomized data column and 1 task column.
Each row in the data column consists of multiple sentences, of which precisely one sentence contains a blank. Your task is to determine the appropriate pronoun to fill in the blank, and enter it in the “Pronoun” column. Here, appropriate means correct in both form and case.
The tasks are designed to be unambiguous, so please provide only one solution and do not reorder the data.
Example:
  • The driver felt unhappy because he did not make enough money. The driver wondered whether ___ should take out a loan.
    => he
Appendix F Experimental Details
We use one 40GB NVIDIA A100 GPU for inference with most models, but we require two GPUs for OPT-30B and four for OPT-66B and the Llama-2-70B base and chat models. We access all models using the Huggingface Transformers library (Wolf et al., 2020) and use the minicons library for pseudo log likelihood evaluation (Misra, 2022).
Appendix G Prompting
Table 3 shows all 10 prompt templates we use to present our task instances to chat models. The first five are open-ended and the last five provide the model with pronoun options to choose from. These templates are used within vanilla and chain-of-thought prompting paradigms as described below.
ID Template
[…]
Table 3: Prompting templates for chat models, where “task” is filled with each dataset instance, and “options” is a list of four pronouns to choose from, all in the correct case.
G.1 Vanilla prompting
With FLAN-T5, vanilla prompting only requires instantiating templates with task instances, whereas Llama-2-chat requires special formatting with INST. Instantiated examples of one template are shown for both models in Table 4. The number of maximum new tokens is set to 5 for FLAN-T5 and 20 for Llama-2-chat based on experimentation.
Model type Example input
FLAN-T5 Fill in the blank with the appropriate pronoun.
The accountant had just eaten a big meal so her stomach
was full. The accountant was asked about ___ charges for
preparing tax returns.
OPTIONS:
- his
- her
- their
- xyr
Llama-2-chat [INST] Fill in the blank with the appropriate pronoun.
The accountant had just eaten a big meal so her stomach
was full. The accountant was asked about ___ charges for
preparing tax returns.
OPTIONS:
- his
- her
- their
- xyr [/INST]
Table 4: Example input using template 8 for FLAN-T5 and Llama-2-chat models.
G.2 Chain-of-thought prompting
We focus on FLAN-T5-xxl for chain-of-thought experiments, and use the strong zero-shot prompts for reasoning proposed by Kojima et al. (2022) (“Let’s think step by step”) and Zhou et al. (2023) (“Let’s work this out in a step by step way to be sure we have the right answer”), which we append after the template. Following their codebases, we first allow the models to generate a chain of thought (with 128 maximum new tokens). Then, we append the chain of thought after the question and elicit the final answer with the string “Therefore, the correct pronoun is,” allowing the model to generate up to 10 new tokens. This two-step process is illustrated with examples in Table 5. We save both the final answer and the chain of thought for later analysis.
Chain-of-thought prompt Example input
Kojima et al. (2022) Q: Fill in the blank with the appropriate pronoun.
(eliciting chain of thought)
The accountant had just eaten a big meal so her stomach
was full. The accountant was asked about ___ charges for
preparing tax returns.
A: Let’s think step by step.
Kojima et al. (2022) Q: Fill in the blank with the appropriate pronoun.
(eliciting final answer)
The accountant had just eaten a big meal so her stomach
was full. The accountant was asked about ___ charges for
preparing tax returns.
A: Let’s think step by step. {generated chain of thought}
Therefore, the correct pronoun is
Zhou et al. (2023) Q: Fill in the blank with the appropriate pronoun.
(eliciting chain of thought)
The accountant had just eaten a big meal so her stomach
was full. The accountant was asked about ___ charges for
preparing tax returns.
A: Let’s work this out in a step by step way to be sure we
have the right answer.
Zhou et al. (2023) Q: Fill in the blank with the appropriate pronoun.
(eliciting final answer)
The accountant had just eaten a big meal so her stomach
was full. The accountant was asked about ___ charges for
preparing tax returns.
A: Let’s work this out in a step by step way to be sure we
have the right answer. {generated chain of thought}
Therefore, the correct pronoun is
Table 5: Example input using template 3 for evaluating FLAN-T5-xxl with two types of chain-of-thought prompting. Prompting happens in two phases regardless of the choice of prompt: eliciting the chain of thought and eliciting the final answer.
he/him/his Refer to caption
she/her/her Refer to caption
they/them/their Refer to caption
xe/xem/xyr Refer to caption
Figure 8: Counts of pronoun predictions from all models, in the absence of context. The random baseline shows counts if each pronoun set was chosen equally often.
Appendix H Context-free pronoun predictions by model
Figure 8 shows per-model pronoun predictions in the absence of context. All models predict he/him/his more frequently than she/her/her, which is in turn predicted more frequently than they/them/their and xe/xem/xyr. However, encoder-only models are more balanced in their predictions across the four pronoun sets, compared to decoder-only models which show very stark differences in pronoun predictions.
Appendix I Results with Vanilla and Chain-of-Thought Prompting
I.1 Vanilla prompting
Prompting is a different model evaluation mechanism than log likelihoods, with higher task demands that lead to lower performance than log likelihoods with both base models and instruction fine-tuned chat models (Hu and Levy, 2023; Hu and Frank, 2024; Kauf et al., 2024). We thus expect vanilla prompting results (using the prompts listed in Appendix G) to be worse than results with log likelihoods. Indeed, Figure 9(a) shows that Llama-2-chat prompting performance is lower than Llama-2 evaluated with log likelihoods, even with no distractors. Figure 9(b) shows the results of standard prompting with FLAN-T5, an encoder-decoder model which shows similar patterns of degradation to decoder-only models. Bigger models are mostly better and degrade more gracefully than the smaller ones, but there remains a lot of variance across prompts, as shown in the box plots.
(a) Llama-2-chat
(b) FLAN-T5
Figure 9: Performance of chat models (Llama-2-chat and FLAN-T5) with additional distractors, using vanilla prompting. The boxplots show the range of performance across 10 different templates.
I.2 Chain-of-thought prompting
As FLAN-T5-xxl shows strong performance with low variance compared to all the other chat models we consider, we focus on this model for additional evaluation with chain-of-thought prompting. Zero-shot chain-of-thought prompting encourages models to think step-by-step, which could in theory produce much better results on pronoun fidelity. While chain-of-thought prompting is excessive for a task as simple as pronoun fidelity, it might encourage the model to explicitly list the referents and associated pronouns, which could help the model predict the correct pronoun with higher accuracy. In practice, however, we find that it leads to worse performance, potentially due to hallucination.
Figure 10 shows the pronoun fidelity of FLAN-T5-xxl with different types of prompting based on the final answer the model provides. Both types of chain-of-thought prompting worsen performance and increase the variance across prompts compared to vanilla prompting.
When examining model-generated answers and chains of thought, we found that FLAN-T5-xxl does not in fact solve the problem step by step as the instruction suggests. Instead, the chain of thought often already contains an answer, and the final answer is not necessarily the same as this one. Therefore, we also plot performance using answers from the model-generated chain of thought in Figure 11. Once again, performance with 1-5 distractors is much lower, showing that chain-of-thought prompting degrades performance compared to vanilla prompting. However, with no distractors, performance is almost exactly the same as vanilla prompting, as models simply generate the answer within the chain of thought. This reinforces that chain-of-thought is unnecessary for a task this simple.
Figure 10: Performance of FLAN-T5-xxl with distractor sentences, comparing vanilla prompting to two types of chain-of-thought prompting. Here, the model’s final answers are used for evaluation and the boxplots show the range of performance across 10 different templates.
Figure 11: Performance of FLAN-T5-xxl with distractor sentences, comparing vanilla prompting to two types of chain-of-thought prompting. Here, the model’s chain of thought is used for evaluation and the boxplots show the range of performance across 10 different templates.